OpenStudy (coolcat):

what is 4y^2-81=0 URGENT PLEASE ANSWER. what is 4y^2-81=0 URGENT PLEASE ANSWER. @Mathematics

OpenStudy (anonymous):

y = 9/2 or -9/2

OpenStudy (coolcat):

thanks! u worked it out quickly! how?

OpenStudy (anonymous):

4y^2=81 y^2= 81/4 y= 9/2; -9/2

OpenStudy (anonymous):

it is simple based on formula \[(a^2 - b^2) = (a-b)(a+b)\]

OpenStudy (anonymous):

in this case a = 2y and b = 9

OpenStudy (anonymous):

\[4y^2=81, \sqrt{4y}=\sqrt{81}, 2y=9, y=9/2\]

OpenStudy (anonymous):

it's a difference of squares. 4y^2 is (2y)^2 and 81 is 9^2 It becomes (2y-9)(2y+9) = 0 set each of the terms equal to zero. 2y-9=0 and 2y+9=0 You get y=9/2 or -9/2